# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@ Appendix C Testing and comparison to CLASS
SymBoltz’ code repository is set up with continuous integration that runs several tests and builds updated documentation pages every time changes to the code are committed. In particular, this compares the solution for $\mathrm{\Lambda CDM}$ with CLASS for many variables solved by the background, thermodynamics and perturbations (using the options write_background, write_thermodynamics and k_output_values). These are the basis for all derived quantities like luminosity distances, matter and CMB power spectra, which are also compared. The checks pass when the quantities agree within a small tolerance. We do not compare directly against more codes like CAMB, but CLASS has already been compared extensively with CAMB with excellent agreement (Lesgourgues, 2011b). The comparison takes a lot of space and is not included here, but is found in the documentation linked from SymBoltz’ repository.
Another test checks that integration of the background and perturbations equations are stable throughout parameter space. As the equations are very stiff and SymBoltz does not rely on approximations for relieving it, one could imagine that the integration would be stable for some parameter values and unstable for others. The test creates a box in parameter space $\pm 50\%$ around a fiducial set of realistic parameter values, draws several sets of parameter values from that space with Latin hypercube sampling (to efficiently cover parameter space) and integrates the background and perturbations for each such set. All parameter samples are found to integrate successfully without warnings and errors.

### v3

@@ Appendix C Testing and comparison to CLASS
SymBoltz’ code repository is set up with continuous integration that runs several tests and builds updated documentation pages every time changes to the code are committed. In particular, this compares the solution for $\mathrm{\Lambda CDM}$ with CLASS for many variables solved by the background, thermodynamics, and perturbations (using the options write_background, write_thermodynamics, and k_output_values). These are the basis for all derived quantities, such as luminosity distances, matter, and CMB power spectra, which are also compared. The checks pass when the quantities agree within a small tolerance. We do not compare directly against additional codes such as CAMB, but CLASS has already been compared extensively with CAMB with excellent agreement (Lesgourgues 2011b). This comparison is found in SymBoltz’ documentation. [^25] https://hersle.github.io/SymBoltz.jl/stable/comparison/
Another test checks that integration of the background and perturbations equations is stable throughout parameter space. As SymBoltz integrates stiff equations without approximations, we could imagine that the integration would be stable for some parameter values and unstable for others. The test creates a box in parameter space $\pm 50\%$ around a fiducial set of realistic parameter values, draws several sets of parameter values from that space with Latin hypercube sampling (to efficiently cover the box), and integrates the background and perturbations for each such set. All samples are found to integrate successfully.